# Valid Stopping for LLM Generation via Empirical Dynamic Formal Lift

### v1

@@ Appendix A Skeleton Details and Diagnostics
A.1 Automated Skeletons and Task-Agnostic Families
@@
Building on this skeleton framework, our approach centers on measuring information lift—the log-likelihood gain when the full model assigns higher probability to tokens than the skeleton expects. This lift quantifies evidence accumulation as generation proceeds, providing the foundation for our stopping criterion.
Definition A.1 (Clipped lift).
@@
Our statistical testing framework operates on the formal null and alternative hypotheses:
$H_{0}:\quad$ $\mathbb{E}\left[\sum_{t=1}^{\tau}X_{t}\mid\text{stop at }\tau\right]\leq\tau_{\text{sufficient}}$ (7)
$H_{1}:\quad$ $\mathbb{E}\left[\sum_{t=1}^{\tau}X_{t}\mid\text{stop at }\tau\right]>\tau_{\text{sufficient}}$ (8)
where $\tau_{\text{sufficient}}$ represents the threshold for information sufficiency. Under $H_{0}$, the cumulative lift remains below the sufficiency threshold, indicating inadequate evidence accumulation. Under $H_{1}$, cumulative lift exceeds the threshold, warranting a stopping decision.
To illustrate this framework, consider a toy example with vocabulary $\{$“unit”, “price”, “\$0.67”}\} and two-step reasoning. The full model $P$ has access to worked examples while skeleton $S$ sees only the instruction. At step 1, both models assign moderate probability to “unit” ($P$: 0.4, $S$: 0.3), yielding lift $X_{1}=\log(0.4/0.3)=0.29$ nats. At step 2, the full model strongly favors the correct answer “\$0.67” ($P$: 0.8) while the skeleton remains uncertain ($S$: 0.2), producing $X_{2}=\log(0.8/0.2)=1.39$ nats. The cumulative lift $\sum X_{t}=1.68$ nats indicates evidence accumulation, and our e-process determines whether this exceeds the statistical boundary for certification.
@@
Appendix B Automated Skeleton Recipes
Table 8 demonstrates that automated skeletons perform comparably to hand-designed alternatives, with TPCA varying by 16 tokens (19% relative variation, from 84.3 to 100.3) across all methods while maintaining consistent empirical risk control (0.083–0.094, all within target $\delta=0.1$).
@@ Appendix B Automated Skeleton Recipes
Validation. Reuse the KL and $\rho(X_{t},H_{t})$ diagnostics (Section 2.2) and accept candidates within 2–10 nats and $\rho<-0.5$.
Appendix C Detailed Deployment Blueprint
@@
C.1 Four-Step Deployment Procedure
@@
C.2 Diagnostic Workflow for Skeleton Construction
Effective skeleton design requires systematic validation: (1) Construct candidate skeleton using domain-appropriate method, (2) Validate on n=50 examples by computing KL(P——S) and $\rho(X_{t},H_{t})$—reject if $KL<2$, $KL>10$, or $\rho>-0.3$, (3) Manual inspection of 20 generated examples to verify that $M_{t}$ peaks align with semantic completion points, (4) Ablate clip bound BB—if many $X_{t}=B$ (saturated), increase $B$; if $X_{t}\ll B$ always, decrease $B$.
Appendix D Comprehensive Failure Mode Analysis
@@
D.1 Visualizing the Correctness-Sufficiency Gap
Figure 4 visualizes the fundamental distinction between information sufficiency (what EDFL certifies) and factual correctness (what safety-critical applications require). The Venn diagram shows that while 72% of sequences achieve both high lift and correctness, 16% exhibit high lift but produce incorrect answers—representing EDFL’s core limitation.
@@ D.1 Visualizing the Correctness-Sufficiency Gap
Figure 4: Information sufficiency vs. factual correctness. EDFL certifies high lift (blue), which correlates with but does not guarantee correctness (green). The 16% gap (average across datasets) represents confident incorrect answers—our method’s fundamental limitation for safety-critical deployment.
@@
[…]
@@
[…]
Our comprehensive failure mode analysis reveals quantified limitations across four categories. Type I errors (early stopping) occur at rates of 6.1-11.4% across datasets, aligning with our target $\delta=0.10$ after conservative inflation. Type II errors (late stopping) remain low at 2.9-6.3%, indicating effective boundary calibration. Certified but incorrect answers represent the most significant limitation, affecting 13.2-22.7% of stopped sequences depending on domain complexity—ProofWriter shows the smallest gap (13.2%) due to structured reasoning, while LegalBench exhibits the largest (22.7%) due to complex argumentation patterns. Drift-induced failures occur in less than 2% of sequences, demonstrating effective adaptive reset mechanisms.
Appendix E Skeleton Construction Details
@@
Prompt compression systematically removes contextual information while preserving task structure. The procedure involves parsing the full prompt into components (instruction, examples, context, query), retaining only the instruction and query while removing examples and context, then constructing $P_{\text{skeleton}}$ as the simplified prompt. For example, in GSM8K, the full prompt “You are a math tutor. Solve step-by-step. Example 1: 3 apples cost \$2, find cost of 7. Solution: Unit price = \$2/3. Then 7*(2/3)=\$4.67. [3 more examples…] Question: {query}” becomes the skeleton “Solve this math problem: {query}”. This approach proves effective for reasoning tasks where examples provide significant guidance but the core task structure remains intact.
@@
Table 11: Skeleton quality validation across datasets using prompt compression method.
Dataset $\mathrm{KL}(P||S)$ $\mathrm{KL}(S||P)$ $\rho(X_{t},H_{t})$
@@
Appendix F Grid Analysis
@@
Table 12: Grid approximation quality (fraction within multiples of optimal $\lambda^{\star}$).
Grid $K$ Within 2× Within 1.5× Avg regret/tok
@@
Appendix G Skip Predictor Validation
@@
Appendix H Sensitivity to Inflation
@@
The aggregate results across all datasets demonstrate clear trends in empirical risk reduction as inflation increases. Without inflation, average empirical risk reaches 0.131 ± 0.018, substantially exceeding our target $\delta=0.1$. Light inflation (1.15×, 1.2×) provides partial improvement to 0.098 ± 0.013, while our default strategy achieves 0.084 ± 0.009, comfortably below the target with reasonable variance. Heavy inflation (1.5×, 2.0×) further reduces risk to 0.069 ± 0.007 but at significant efficiency cost, making the default strategy optimal for practical deployment.
Table 14: Empirical risk vs. inflation factors (mean $\pm$ std across datasets).
@@
Per-dataset analysis confirms consistent benefits of default inflation across diverse domains, with all datasets achieving empirical risk below the 0.1 target when inflation is applied, compared to systematic violations without inflation. The effectiveness proves particularly pronounced for challenging datasets like ASQA and LegalBench, where complex reasoning patterns benefit most from conservative parameter estimation.
@@
These results demonstrate that default inflation provides robust risk control across all datasets, with empirical risk consistently below target levels compared to systematic violations without inflation.
Appendix I Expanded Experimental Results
@@
The gate reduces the HighLift+Incorrect rate by 41% (from 18.5% to 10.9%) with modest TPCA increase (+7.1 tokens) and overhead growth (+4%, Table 17), narrowing the sufficiency–correctness gap while preserving anytime-valid guarantees. The 83% verification reduction is computed as follows: the gate successfully validates 83% of EDFL stopping decisions (Table 17), requiring full verification only for the 17% of stops that fail the gate plus any sequences that continue past EDFL’s threshold.
Table 17: Hybrid gate effect (GSM8K + HotpotQA + ASQA, $n=500$ each, $\delta=0.1$).
Setting Acc HighLift+Incorrect TPCA Ovhd
EDFL (base) 76.8% 18.5% 90.7 12%
+ Sentence boundary only 79.6% 14.7% 93.2 13%
+ Boundary + verifier 83.4% 10.9% 97.8 16%
@@
Appendix J Proofs
Proof.
The proof of Lemma 3.2 establishes that optional skipping preserves the supermartingale property essential for statistical validity. Because our gate only waits (never triggers sooner), it cannot inflate Type-I error; it only trades a few tokens for better correctness. Formally, optional stopping for nonnegative supermartingales plus the fact that the gate only delays stopping (never advancing) implies validity is unchanged. See, e.g., time-uniform concentration via e-processes (Howard et al., 2021a; Vovk and Wang, 2021b; Waudby-Smith and Ramdas, 2023). When $\hat{s}_{t}=1$ triggers skipping, we set $X_{t}=0$, yielding $\tilde{Z}_{t}=0-\hat{\mu}_{t-1}\leq 0$ since $\hat{\mu}_{t-1}\geq 0$ (all $X_{i}\geq 0$). For $\lambda>0$ and $\tilde{Z}_{t}\leq 0$, the conditional expectation $\mathbb{E}[\exp(\lambda\tilde{Z}_{t})|\mathcal{F}_{t-1}]=\exp(\lambda\tilde{Z}_{t})\leq 1\leq\exp\left(\frac{\lambda^{2}(\hat{v}_{t-1}+\eta)}{2(1-c\lambda)}\right)$ remains bounded by our penalty function. The e-process update $\mathbb{E}[M_{t}|\mathcal{F}_{t-1}]=M_{t-1}\cdot\mathbb{E}[\exp(\lambda\tilde{Z}_{t}-\psi_{t}(\lambda))|\mathcal{F}_{t-1}]\leq M_{t-1}$ preserves the supermartingale property, and by optional stopping theorem (since $\hat{s}_{t}\in\mathcal{F}_{t-1}$), Theorem 3.1 continues to hold. ∎
The proof of Theorem 3.3 extends validity to adaptive segment budgeting by defining segment-wise e-processes $M^{(j)}_{t}=\sum_{k}w_{k}\exp\left(\lambda_{k}\sum_{i\in[t_{j},t]}\tilde{Z}_{i}-\sum_{i\in[t_{j},t]}\psi_{i}(\lambda_{k})\right)$ where each segment $j$ satisfies $\mathbb{P}\left(\sup_{t}M^{(j)}_{t}\geq 1/\delta_{j}\right)\leq\delta_{j}$ by Theorem 3.1. Within segment $j$, apply Theorem 3.1 with budget $\delta_{j}$. Union bound over segments yields $\sum_{j}\delta_{j}=\delta$. Since $\delta_{j}$ decays as $j^{-2}$, the series converges. This is a standard convergent-series budget allocation; cf. segment-wise testing with e-processes (Howard et al., 2021a; Ramdas et al., 2023).The union bound over segments yields $\mathbb{P}\left(\exists j,t:M^{(j)}_{t}\geq 1/\delta_{j}\right)\leq\sum_{j=1}^{\infty}\delta_{j}=\sum_{j=1}^{\infty}\frac{6\delta}{\pi^{2}j^{2}}=\delta$, completing the proof that adaptive resets maintain overall error control at level $\delta$.
Appendix K Computational Overhead Breakdown
Detailed profiling of Sequential-EDFL reveals that the 12.1% computational overhead stems primarily from skeleton logit computation, which dominates at 6.5% of baseline generation time. The remaining components contribute modestly: e-process computation across 12 $\lambda$ values adds 2.1%, drift estimation contributes 1.2%, information lift calculation requires 1.0%, entropy-slope computation adds 0.8%, and EMA updates consume 0.6%. This breakdown suggests optimization opportunities through skeleton caching or distilled models for the dominant component.
Component Time (ms) % of baseline
Baseline generation 520 $\pm$ 18
Skeleton logits 34 $\pm$ 4 6.5%
Lift $X_{t}$ 5 $\pm$ 1 1.0%
EMA updates ($\hat{\mu},\hat{v}$) 3 $\pm$ 1 0.6%
[…]
Entropy-slope 4 $\pm$ 1 0.8%
Drift estimator 6 $\pm$ 1 1.2%
Total overhead 63 $\pm$ 6 12.1%
Appendix L Error Categorization
@@
Dataset Correct Early (FP) Late (FN) Timeout
GSM8K 87.9% 7.3% 3.2% 1.6%
HotpotQA 82.7% 9.4% 5.6% 2.3%
ASQA 81.2% 10.3% 6.1% 2.4%
TruthfulQA 85.8% 8.1% 4.4% 1.7%
ProofWriter 89.6% 6.1% 2.9% 1.4%
LegalBench 79.9% 11.4% 6.3% 2.4%
Appendix M Reproducibility Checklist
Appendix N Algorithmic Specification
@@
Appendix O Extended Related Work
@@
Sequential-EDFL builds upon sequential testing theory, particularly e-processes Vovk and Wang (2021a); Ramdas et al. (2023); Howard et al. (2021a) that provide anytime-valid inference with roots in Ville’s martingale theory Ville (1939) and modern confidence sequences Robbins (1970); Lai (1976); Howard et al. (2021b). Self-normalized e-processes Howard et al. (2021a); Waudby-Smith and Ramdas (2023) handle unknown centering via running variance estimation, directly inspiring our empirical-Bernstein extension. Recent work on sequential inference in LLMs Lew et al. (2023) demonstrates the value of adaptive stopping, though without anytime-valid guarantees. Safe testing Grünwald et al. (2019) and betting scores Waudby-Smith and Ramdas (2021) provide additional theoretical grounding and baseline comparisons for our anytime-valid framework.
@@
Conformal prediction represents a complementary approach to uncertainty quantification that offers finite-sample guarantees through split conformal methods Vovk et al. (2005); Angelopoulos and Bates (2021), though requiring calibration sets. Extensions encompass conformalized quantile regression Romano et al. (2019), distribution-free inference Lei et al. (2018), and covariate shift adaptation Tibshirani et al. (2019), with online variants Gibbs and Candes (2021) and time series extensions Darrin and Tibshirani (2023) addressing distribution shift. However, these approaches don’t naturally accommodate token-level stopping decisions in sequential generation, motivating our e-process framework.
@@
Within the LLM uncertainty landscape, existing approaches include entropy-based methods Malinin and Gales (2018), semantic consistency checks Manakul et al. (2023), verbalized confidence Kadavath et al. (2022), and hallucination detection Varshney et al. (2023); Fadeeva et al. (2024), yet these lack the formal statistical guarantees that our framework provides through frequentist control with anytime-valid error rates. Risk-aware decoding encompasses constrained generation Hokamp and Liu (2017), calibration techniques Guo et al. (2017); Minderer et al. (2021); Kumar et al. (2019), and selective abstention Geifman and El-Yaniv (2017) using static thresholds, alongside out-of-distribution detection Hendrycks and Gimpel (2017); Liang et al. (2018); Lee et al. (2018) for uncertainty estimation. Our work extends this paradigm by providing sequential, anytime-valid control that dynamically adapts to evidence accumulation patterns.

### v2

@@ Appendix B Skeleton Details and Diagnostics
B.1 Automated Skeletons and Task-Agnostic Families
@@
Building on this skeleton framework, our approach centers on measuring information lift, which is the log-likelihood gain when the full model assigns higher probability to tokens than the skeleton expects. This lift quantifies evidence accumulation as generation proceeds, providing the foundation for our stopping criterion.
Definition B.1 (Clipped lift).
@@
Our statistical testing framework defines the error event as stopping prematurely relative to information sufficiency. Formally, let $S_{t}$ be the sufficiency indicator at time $t$ (as defined in Section 4: $S_{t}=1$ if the answer at time $t$ matches the final answer $A_{T}$ and the sequence is syntactically complete). The stopping rule $\tau=\inf\{t:M_{t}\geq 1/\delta\}$ controls the probability:
$\mathbb{P}\left(\text{stop at time }\tau\text{ with }S_{\tau}=0\right)\leq\delta.$ (7)
This guarantee means that with probability at least $1-\delta$, we do not stop before the sequence becomes sufficient (answer-stable and complete). The e-process $M_{t}$ accumulates evidence through information lift $X_{t}$, and the threshold $1/\delta$ provides the statistical boundary ensuring this error control regardless of stopping time (anytime-validity).
To illustrate this framework, consider a toy example with vocabulary $\{$“unit", “price", “\$0.67"}\} and two-step reasoning. The full model $P$ has access to worked examples while skeleton $S$ sees only the instruction. At step 1, both models assign moderate probability to “unit" ($P$: 0.4, $S$: 0.3), yielding lift $X_{1}=\log(0.4/0.3)=0.29$ nats. At step 2, the full model strongly favors the correct answer “\$0.67" ($P$: 0.8) while the skeleton remains uncertain ($S$: 0.2), producing $X_{2}=\log(0.8/0.2)=1.39$ nats. The cumulative lift $\sum X_{t}=1.68$ nats indicates evidence accumulation, and our e-process determines whether this exceeds the statistical boundary for certification.
@@
Appendix C Automated Skeleton Recipes
Table 9 demonstrates that automated skeletons perform comparably to hand-designed alternatives, with TPCA varying by 16 tokens (19% relative variation, from 84.3 to 100.3) across all methods while maintaining consistent premature-stop rate control (0.083–0.094, all within target $\delta=0.1$).
@@ Appendix C Automated Skeleton Recipes
Validation. Reuse the KL and $\rho(X_{t},H_{t})$ diagnostics and accept candidates within 2–10 nats and $\rho<-0.5$.
Appendix D Detailed Deployment Blueprint
@@
D.1 Four-Step Deployment Procedure
@@
D.2 Diagnostic Workflow for Skeleton Construction
Effective skeleton design requires systematic validation. First, construct candidate skeleton using domain-appropriate method. Second, validate on n=50 examples by computing KL(P||S) and $\rho(X_{t},H_{t})$; reject if $KL<2$, $KL>10$, or $\rho>-0.3$. Third, manually inspect 20 generated examples to verify that $M_{t}$ peaks align with semantic completion points. Fourth, ablate clip bound $B$: if many $X_{t}=B$ (saturated), increase $B$; if $X_{t}\ll B$ always, decrease $B$.
Appendix E Comprehensive Failure Mode Analysis
@@
E.1 Visualizing the Correctness-Sufficiency Gap
Figure 4 visualizes the fundamental distinction between information sufficiency (what EDFL certifies) and factual correctness (what safety-critical applications require). The Venn diagram shows that while 72% of sequences achieve both high lift and correctness, 16% exhibit high lift but produce incorrect answers, representing EDFL’s core limitation.
@@ E.1 Visualizing the Correctness-Sufficiency Gap
Figure 4: Information sufficiency vs. factual correctness. EDFL certifies high lift (blue), which correlates with but does not guarantee correctness (green). The 16% gap (average across datasets) represents confident incorrect answers, which is our method’s fundamental limitation for safety-critical deployment.
@@
After numerical expression 68% 8.2% “…= \$4.67" $\leftarrow$ STOP
After conclusion marker 21% 11.7% “Therefore the answer is 7" $\leftarrow$ STOP
@@
After punctuation 4% 15.0% “…total cost is \$4.67." $\leftarrow$ STOP
Our comprehensive failure mode analysis reveals quantified limitations across four categories. Type I errors (early stopping) occur at rates of 6.1-11.4% across datasets, aligning with our target $\delta=0.10$ after conservative inflation. Type II errors (late stopping) remain low at 2.9-6.3%, indicating effective boundary calibration. Certified but incorrect answers represent the most significant limitation, affecting 13.2-22.7% of stopped sequences depending on domain complexity. ProofWriter shows the smallest gap (13.2%) due to structured reasoning, while LegalBench exhibits the largest (22.7%) due to complex argumentation patterns. Drift-induced failures occur in less than 2% of sequences, demonstrating effective adaptive reset mechanisms.
Appendix F Skeleton Construction Details
@@
Prompt compression systematically removes contextual information while preserving task structure. The procedure involves parsing the full prompt into components (instruction, examples, context, query), retaining only the instruction and query while removing examples and context, then constructing $P_{\text{skeleton}}$ as the simplified prompt. For example, in GSM8K, the full prompt “You are a math tutor. Solve step-by-step. Example 1: 3 apples cost \$2, find cost of 7. Solution: Unit price = \$2/3. Then 7*(2/3)=\$4.67. [3 more examples…] Question: {query}" becomes the skeleton “Solve this math problem: {query}". This approach proves effective for reasoning tasks where examples provide significant guidance but the core task structure remains intact.
@@
Table 12: Skeleton quality validation across datasets (prompt compression).
Dataset $\mathrm{KL}(P\|S)$ $\mathrm{KL}(S\|P)$ $\rho(X_{t},H_{t})$
@@
Appendix G Grid Analysis
@@
Table 13: Grid approximation quality (within multiples of $\lambda^{\star}$).
[…]
@@
Appendix H Skip Predictor Validation
@@
Appendix I Sensitivity to Inflation
@@
The aggregate results across all datasets demonstrate clear trends in premature-stop rate reduction as inflation increases. Without inflation, average premature-stop rate reaches 0.131 ± 0.018, substantially exceeding our target $\delta=0.1$. Light inflation (1.15×, 1.2×) provides partial improvement to 0.098 ± 0.013, while our default strategy achieves 0.084 ± 0.009, comfortably below the target with reasonable variance. Heavy inflation (1.5×, 2.0×) further reduces premature-stop rate to 0.069 ± 0.007 but at significant efficiency cost, making the default strategy optimal for practical deployment.
Table 15: Premature-stop rate vs. inflation factors (mean $\pm$ std across datasets).
@@
Per-dataset analysis confirms consistent benefits of default inflation across diverse domains, with all datasets achieving premature-stop rate below the 0.1 target when inflation is applied, compared to systematic violations without inflation. The effectiveness proves particularly pronounced for challenging datasets like ASQA and LegalBench, where complex reasoning patterns benefit most from conservative parameter estimation.
@@
These results demonstrate that default inflation provides robust risk control across all datasets, with premature-stop rate consistently below target levels compared to systematic violations without inflation.
Appendix J Detailed Experimental Analysis
This section provides comprehensive analysis of stopping patterns, failure modes, sensitivity analysis, skeleton comparisons, ablation studies, human evaluation, and qualitative case studies. These details support the main experimental findings reported in Section 4.
J.1 Stopping Patterns and Failure Mode Analysis
Token-level stopping patterns: Analysis of 200 stopped sequences from GSM8K reveals systematic patterns in where the method chooses to stop, with significant implications for answer quality. The 5× higher error rate for mid-sentence stops (42.3% vs. 8.2%) suggests that syntactic completeness provides a crucial signal orthogonal to information lift. Post-hoc filtering to allow stops only at sentence boundaries reduces overall error rates from 8.3% to 6.1% on GSM8K, demonstrating how linguistic constraints can enhance statistical stopping criteria. Quantitative breakdown appears in Table 11 (Appendix E).
Failure mode taxonomy: Detailed analysis of the High Lift + Incorrect sequences reveals four distinct failure patterns with different underlying causes and potential mitigations (Table 17).
Failure Type % of Errors Lift Pattern Mitigation Strategy
Factual hallucination 28% Sudden spike Knowledge base lookup
J.2 Sensitivity Analysis
[…]
J.3 Skeleton Comparison Analysis
J.4 Hyperparameter Ablation Analysis
The ablation results in Table 7 (main paper) provide clear guidance for hyperparameter selection: $K=12$ optimally balances efficiency and computational overhead, $\eta=\alpha$ maintains premature-stop rate near the target $\delta$, $\alpha=0.2$ provides robust performance across diverse settings, and $\tau_{d}=0.10$ optimizes the risk-efficiency trade-off. These findings demonstrate that our method’s performance remains stable across reasonable parameter ranges, indicating robustness to hyperparameter choices.
J.5 Human Evaluation and Case Studies
The human evaluation results in Table 18 reveal that early stops sacrifice slight completeness (0.4 points) while maintaining similar correctness, validating that information sufficiency correlates with, but doesn’t guarantee, factual correctness. This finding supports our theoretical framework while highlighting the distinction between statistical confidence and semantic quality, consistent with prior work on confidence-correctness gaps tian2023just; kadavath2022language.
Metric Sequential-EDFL Fixed (150 tok)
Completeness 3.8 $\pm$ 0.9 4.2 $\pm$ 0.7
Correctness 3.6 $\pm$ 1.1 3.7 $\pm$ 1.0
Three qualitative case studies illuminate Sequential-EDFL’s behavioral patterns across different scenarios. A successful multi-hop RAG example on HotpotQA demonstrates the method’s ability to track evidence accumulation: for “What is the population of the metro area where Tesla’s headquarters is located?", the system shows moderate lift (2-4 nats) while identifying “Austin" in tokens 1-20, then experiences lift spikes (8-12 nats) when extracting “2.3 million" in tokens 21-43, triggering stopping at token 43 when $M_{t}=17.8>u_{1}=16.4$ with the correct answer. The context ablation skeleton (no retrieval) remains uncertain about both location and population, yielding large lift when evidence resolves both aspects.
Conversely, a TruthfulQA failure case demonstrates the critical distinction between confidence and correctness: for “What happens if you crack your knuckles too much?", Sequential-EDFL stopped at token 38 with the incorrect answer “Cracking your knuckles too much will cause arthritis and joint damage over time" (gold: “Nothing harmful happens"). The model confidently asserted this misconception with high $X_{t}$ (6-10 nats) because the full model strongly believed this claim while the skeleton remained uncertain, illustrating that information lift certifies confidence rather than correctness. Finally, the adaptive reset mechanism’s importance emerges in an ASQA answer about the Industrial Revolution (2847 tokens total), where three resets at tokens 873, 1692, and 2401 enabled comprehensive coverage as topics shifted from economic to technological to social to global impacts. Human raters scored 4.2/5 completeness with resets versus 2.1/5 without, confirming the mechanism’s effectiveness for long-form generation.
Appendix K Expanded Experimental Results
K.1 Open-Ended Generation Tasks
To assess applicability beyond single-turn QA, we evaluate EDFL on two open-ended generation tasks: dialogue (Wizard-of-Wikipedia dinan2019wizard) and abstractive summarization (CNN/DailyMail hermann2015teaching). For dialogue, the skeleton removes Wikipedia grounding (context ablation); for summarization, it uses temperature scaling ($\tau=1.8$) as no external information source exists. Since these tasks lack ground-truth “correctness," we measure fluency (perplexity of continuations), informativeness (ROUGE-L vs reference), and human preference (50 examples, 3 raters, pairwise comparison vs fixed-length baseline).
Results show EDFL reduces tokens by 21-31% on open-ended tasks while maintaining acceptable quality: perplexity increases by 5-7% (acceptable fluency degradation), ROUGE-L drops by 3% (modest informativeness loss), and humans prefer EDFL outputs 54-58% of the time (near-parity with fixed-length). The larger variance ($\pm$6-9 tokens vs $\pm$2-3 on QA) reflects diverse response lengths in open-ended generation. These results suggest EDFL generalizes beyond factual QA, though the information-sufficiency criterion becomes less well-defined without clear grounding sources. For truly open-ended creative tasks (story generation, brainstorming), where no skeleton baseline exists, our method’s applicability remains limited, which is an important scope constraint.
@@
The gate reduces the HighLift+Incorrect rate by 41% (from 18.5% to 10.9%) with modest TPCA increase (+7.1 tokens) and overhead growth (+4%, Table 3), narrowing the sufficiency–correctness gap while preserving anytime-valid guarantees. Under a verification policy that checks all outputs, the gate reduces verifier-call rate from 100% to $\sim$17% (gate validates 83% of stops, requiring full verification only for remaining 17% plus non-stopped sequences), representing a practical systems benefit but not a formal guarantee.
Appendix L Proofs
Proof.
The proof of Lemma 3.2 establishes that optional skipping preserves the supermartingale property essential for statistical validity. Because our gate only waits (never triggers sooner), it cannot inflate Type-I error; it only trades a few tokens for better correctness. Formally, optional stopping for nonnegative supermartingales plus the fact that the gate only delays stopping (never advancing) implies validity is unchanged. See, e.g., time-uniform concentration via e-processes (howard2021timeuniform; vovk2021evalue; waudbysmith2023empirical). When $\hat{s}_{t}=1$ triggers skipping, we set $X_{t}=0$, yielding $\tilde{Z}_{t}=0-\hat{\mu}_{t-1}\leq 0$ since $\hat{\mu}_{t-1}\geq 0$ (all $X_{i}\geq 0$). For $\lambda>0$ and $\tilde{Z}_{t}\leq 0$, the conditional expectation $\mathbb{E}[\exp(\lambda\tilde{Z}_{t})|\mathcal{F}_{t-1}]=\exp(\lambda\tilde{Z}_{t})\leq 1\leq\exp\left(\frac{\lambda^{2}(\hat{v}_{t-1}+\eta)}{2(1-c\lambda)}\right)$ remains bounded by our penalty function. The e-process update $\mathbb{E}[M_{t}|\mathcal{F}_{t-1}]=M_{t-1}\cdot\mathbb{E}[\exp(\lambda\tilde{Z}_{t}-\psi_{t}(\lambda))|\mathcal{F}_{t-1}]\leq M_{t-1}$ preserves the supermartingale property, and by optional stopping theorem (since $\hat{s}_{t}\in\mathcal{F}_{t-1}$), Theorem 3.1 continues to hold. ∎
@@
The proof of Theorem 3.3 extends validity to adaptive segment budgeting by defining segment-wise e-processes $M^{(j)}_{t}=\sum_{k}w_{k}\exp\left(\lambda_{k}\sum_{i\in[t_{j},t]}\tilde{Z}_{i}-\sum_{i\in[t_{j},t]}\psi_{i}(\lambda_{k})\right)$ where each segment $j$ satisfies $\mathbb{P}\left(\sup_{t}M^{(j)}_{t}\geq 1/\delta_{j}\right)\leq\delta_{j}$ by Theorem 3.1. Within segment $j$, apply Theorem˜3.1 with budget $\delta_{j}$. Union bound over segments yields $\sum_{j}\delta_{j}=\delta$. Since $\delta_{j}$ decays as $j^{-2}$, the series converges. This is a standard convergent-series budget allocation; cf. segment-wise testing with e-processes (howard2021timeuniform; ramdas2023gametheoretic).The union bound over segments yields $\mathbb{P}\left(\exists j,t:M^{(j)}_{t}\geq 1/\delta_{j}\right)\leq\sum_{j=1}^{\infty}\delta_{j}=\sum_{j=1}^{\infty}\frac{6\delta}{\pi^{2}j^{2}}=\delta$, completing the proof that adaptive resets maintain overall error control at level $\delta$.
Appendix M Computational Overhead Breakdown
Detailed profiling of Sequential-EDFL reveals that the 12.1% computational overhead stems primarily from skeleton logit computation, which dominates at 6.5% of baseline generation time. The remaining components contribute modestly: e-process computation across 12 $\lambda$ values adds 2.1%, drift estimation contributes 1.2%, information lift calculation requires 1.0%, entropy-slope computation adds 0.8%, and EMA updates consume 0.6%. This breakdown suggests optimization opportunities through skeleton caching or distilled models for the dominant component. Per-token latency analysis for streaming deployments appears in Appendix W.
Appendix N Error Categorization
@@
Appendix O Algorithmic Specification
@@
Component Time (ms) % of baseline
Baseline generation 520 $\pm$ 18
Skeleton logits 34 $\pm$ 4 6.5%
Lift $X_{t}$ 5 $\pm$ 1 1.0%
EMA updates ($\hat{\mu},\hat{v}$) 3 $\pm$ 1 0.6%
[…]
Entropy-slope 4 $\pm$ 1 0.8%
Drift estimator 6 $\pm$ 1 1.2%
Total overhead 63 $\pm$ 6 12.1%
Appendix P Reproducibility Checklist
Appendix Q Extended Related Work
@@
Sequential-EDFL builds upon sequential testing theory, particularly e-processes vovk2021combining; ramdas2023gametheoretic; howard2021timeuniform that provide anytime-valid inference with roots in Ville’s martingale theory ville1939etude and modern confidence sequences robbins1970statistical; lai1976confidence; howard2021time. Self-normalized e-processes howard2021timeuniform; waudbysmith2023empirical handle unknown centering via running variance estimation, directly inspiring our empirical-Bernstein extension. Recent work on sequential inference in LLMs lew2023sequential demonstrates the value of adaptive stopping, though without anytime-valid guarantees. Safe testing grünwald2019safe and betting scores waudbysmith2021estimating provide additional theoretical grounding and baseline comparisons for our anytime-valid framework.
@@
Conformal prediction represents a complementary approach to uncertainty quantification that offers finite-sample guarantees through split conformal methods vovk2005algorithmic; angelopoulos2021gentle, though requiring calibration sets. Extensions encompass conformalized quantile regression romano2019conformalized, distribution-free inference lei2018distribution, and covariate shift adaptation tibshirani2019conformal, with online variants gibbs2021adaptive and time series extensions chen2020conformal addressing distribution shift. However, these approaches don’t naturally accommodate token-level stopping decisions in sequential generation, motivating our e-process framework.
Dataset Correct Early (FP) Late (FN) Timeout
GSM8K 87.9% 7.3% 3.2% 1.6%
HotpotQA 82.7% 9.4% 5.6% 2.3%
ASQA 81.2% 10.3% 6.1% 2.4%
TruthfulQA 85.8% 8.1% 4.4% 1.7%
ProofWriter 89.6% 6.1% 2.9% 1.4%
LegalBench 79.9% 11.4% 6.3% 2.4%
@@
Within the LLM uncertainty landscape, existing approaches include entropy-based methods malinin2018predictive, semantic consistency checks manakul2023selfcheckgpt, verbalized confidence kadavath2022language, and hallucination detection varshney2023stitch; band2024linguistic, yet these lack the formal statistical guarantees that our framework provides through frequentist control with anytime-valid error rates. Risk-aware decoding encompasses constrained generation hokamp2017lexically, calibration techniques guo2017calibration; minderer2021revisiting; kumar2019verified, and selective abstention geifman2017selective using static thresholds, alongside out-of-distribution detection hendrycks2017baseline; liang2017enhancing; lee2018simple for uncertainty estimation. Our work extends this paradigm by providing sequential, anytime-valid control that dynamically adapts to evidence accumulation patterns.